OpenStudy (anonymous):

X-4y=14 in slope intercept form

OpenStudy (tkhunny):

Can you solve for y?

OpenStudy (anonymous):

since the slope int form is y=mx+b subtract x from both sides -4y=-x+14 Isolate y you should be able to do the rest

OpenStudy (anonymous):

I got y=-1/4x+ -3.5? Is that correct?

OpenStudy (anonymous):

@supraron

OpenStudy (anonymous):

No. if you divide by a negative then 1/4 should be positive. and for -3.5 that's good but put in into a fraction it is better so y=1/4x-7/2

OpenStudy (tkhunny):

Missing a sign on x. Give it another go. This is why I would recommend a slightly different methodology. Given this x-4y=14 Add 4y, rather than subtract x. Don't get all married to keeping y on the left. x = 4y + 14 After this, there just are not that many sign problems.

OpenStudy (anonymous):

Got it @Maymayk :) ?

OpenStudy (anonymous):

Kinda can you help me graph it? @supraron

OpenStudy (anonymous):

Sure, since b, which is 7/2 is the y-intercept then one point would be (0/7/2). Now since your slope is 1/4 count 1 space up and 4 spaces to the right from the point (0,7/2) do that several times and draw your line . DONE ^^

OpenStudy (anonymous):

Got it now ?